OpenStudy (anonymous):

find the limit of the sequence lim (-1)^n/2√n n→∞ How come for when the absolute value of this sequence is taken, that the exponent n on the numerator goes away?

OpenStudy (wmj259):

\[\lim_{n \rightarrow \infty}((\frac{ ((-1)^{n} }{ 2\sqrt{n} } \] Is this the correct form?